Solve the equation for the time in flight and then
Answer #1

Given equation is -120-300 sin 25-0.5 × 9.8ť -120 = 300 × 0.422-4.9t- -120 = 126.6-4.9ŕ 4.9r-126.6-120 = 0 t 26.75s